OpenStudy (anonymous):

40y+9 /20 how do i reduce it

OpenStudy (anonymous):

\[\frac{40y+9}{20}\]

OpenStudy (anonymous):

40y/20+9/20 2y=-9/20 y=-9/40

OpenStudy (anonymous):

lol im bad at fractions, but okay i see what ur doing ur splitting them up 40y/20 + 9/20 thats all i get

OpenStudy (anonymous):

yes so your answer is 2y+9/20

OpenStudy (anonymous):

ok okay.. but hows it negative

OpenStudy (anonymous):

9/20 /2 = 9/40

OpenStudy (anonymous):

once you separate you will need to solve for y, so you subtract 9/20 which gives you 2y=-9/20 then divide by 2 y=-9/40